status reports 18-November-2017 09:37:30

No. 266 - Continuation of quadrature operations, temporary star tracker outage, start of the sixteenth occultation season, and entire thirteenth Atmospheric Drag Experiment Campaign

08 April 2014

Report for the period 13 October to 9 November 2013

This reporting period covers four weeks of Venus Express operations. It includes the continuation of quadrature operations, a temporary star tracker outage, start of the sixteenth occultation season, continuation of the first sulphur dioxide campaign, and the entire thirteenth Atmospheric Drag Experiment Campaign.

Cebreros ground station

All Cebreros ground station activities were nominal during this reporting period.

As planned, Cebreros communication passes on 14, 16, 18, 21 and 23 October were skipped to allow sufficient time for the spacecraft to cool between Earth pointings during the quadrature operations.

New Norcia ground station

The Intermediate Frequency and Modem System 1 (IFMS1) of the New Norcia ground station experienced issues with locking [with the spacecraft] during occultation observations since the beginning of the occultation season on 16 October, until 28 October. This problem was investigated by ESTRACK but the underlying reason was not found; the problem did not recur again during this reporting period. While the anomaly was under investigation, NASA Deep Space Network support was requested during occultation ingress and egress, which is when the anomaly is seen to occur.

Problems with the ground station configuration affected the occultation passes on 8 November.

High-accuracy spacecraft ranging

ESA's Cebreros and New Norcia ground stations were used for a Delta Differential One-way Ranging (delta-DOR, or DDOR) measurement at the start of the Cebreros communications pass on 17 October.

These high accuracy ranging measurements are carried out with the Venus Express spacecraft on a regular basis to support the accurate determination of the ephemeris for the planet Venus that is maintained by NASA's Solar System Dynamics Group.

For more information about DDOR, see "Delta-DOR measurements".

Continuation of quadrature operations

The spacecraft continued to operate in the quadrature configuration during this reporting period.

As per routine quadrature operations procedures, the secondary High Gain Antenna, HGA-2, was swapped with the primary, HGA-1, on 29 October to temporarily become the primary means of communication. This occurred while the spacecraft was outside visibility and hence safing commands were uploaded simultaneously. The swap proceeded smoothly and the safing commands were later deleted during the communications pass on the same day. The illumination conditions during quadrature make Earth pointings 'hot' during communications periods with Earth. The spacecraft was kept in the cool orientation for a few hours to recover from these hot pointings, allowing no solar illumination on thermally sensitive faces, and Earth communication passes were skipped every second day during this quadrature season. The telemetry rate was lowered to 38 kbps after the swap.

For more information on quadrature operations see 'Quadrature operations on Venus Express'.

Start of the sixteenth occultation season

The sixteenth occultation season started on 25 October and continued throughout this reporting period.

The first occultation of the season, on 25 October, was a tangential occultation: the spacecraft did not actually pass behind Venus from the point of view of the ground station. Thus the start and end times for the first occultation observation were not recorded in the relevant file at first. The file was later modified to reflect the change due to the tangential occultation.

For more information, see 'Venus Express Earth occultation seasons'.

Temporary star tracker outage

The spacecraft's star tracker experienced an outage from 18:12 UTC to 22:31 UTC on 28 October. An analysis of spacecraft data was carried out, which ruled out outages due to nominal blinding. Earth-orbiting satellites detected an X1-class solar flare on 28 October at 02:03 UTC and the GOES satellite also showed a proton flux increase at the time of the anomaly. Thus it was determined that the cause of the outage was increased solar activity.

The star tracker has experienced no problems since and is operating nominally.

The thirteenth Atmospheric Drag Experiment campaign

The thirteen Atmospheric Drag Experiment (ADE) Campaign started on 30 October, and ended on 9 November, the last day of this reporting period.

Read more about ADE campaigns here.

Summary of main activities

The table below shows a chronology of the main spacecraft bus activities in the reporting period:

Main activities during reporting period
ADE = Atmospheric Drag Experiment; CEB = Cebreros; DDOR = Delta Differential One-way Ranging; DOY = Day of year; MET = Mission elapsed time; NNO = New Norcia; TM = telemetry
MET
(Day)
Date DOY Main Activity
2896 13-Oct-2013 286 CEB communication pass.
2897 14-Oct-2013 287 CEB skipped pass.
2898 15-Oct-2013 288 CEB communication pass.
2899 16-Oct-2013 289 CEB skipped pass.
2900 17-Oct-2013 290 CEB communication pass; DDOR with NNO and CEB.
2901 18-Oct-2013 291 CEB skipped pass.
2902 19-Oct-2013 292 CEB communication pass.
2903 20-Oct-2013 293 CEB communication pass.
2904 21-Oct-2013 294 CEB skipped pass.
2905 22-Oct-2013 295 CEB communication pass.
2906 23-Oct-2013 296 CEB skipped pass; NNO radio science pass.
2907 24-Oct-2013 297 CEB communication pass; problem with primary CEB telemetry chain, VIRTIS and MAG data lost.
2908 25-Oct-2013 298 Start of the sixteenth occultation season; NNO radio science pass; CEB communication pass.
2909 26-Oct-2013 299 CEB communication pass.
2910 27-Oct-2013 300 NNO radio science pass; CEB communication pass.
2911 28-Oct-2013 301 NNO radio science pass; CEB communication pass; Retrieval of VIRTIS and MAG data that were missed due to a problem with the primary TM chain at CEB on 24 October.
2912 29-Oct-2013 302 NNO radio science pass; CEB communication pass; HGA-1 to HGA-2 swap as part of the inbound quadrature operations, Telemetry bit rate change from 228 kbps to 38 kbps.
2913 30-Oct-2013 303 NNO radio science pass; CEB communication pass; start of the thirteenth ADE campaign.
2914 31-Oct-2013 304 NNO radio science pass; CEB communication pass.
2915 1-Nov-2013 305 NNO radio science pass; CEB communication pass.
2916 2-Nov-2013 306 NNO radio science pass; CEB communication pass.
2917 3-Nov-2013 307 NNO radio science pass; CEB communication pass.
2918 4-Nov-2013 308 NNO radio science pass; CEB communication pass.
2919 5-Nov-2013 309 NNO radio science pass; CEB maintenance followed by shortened communication pass.
2920 6-Nov-2013 310 NNO radio science pass; CEB maintenance followed by shortened communication pass.
2921 7-Nov-2013 311 NNO radio science pass; CEB communication pass.
2922 8-Nov-2013 312 NNO radio science pass – only partial data recording due to misconfiguration of station; CEB communication pass.
2923 9-Nov-2013 313 NNO radio science pass; CEB communication pass; end of the sixteenth occultation season.

At the end of the reporting period on 9 November Venus Express was 92.2 million kilometres from Earth. The one-way signal travel time was 304 seconds. The final oxidizer mass was 15.005 kg and the final fuel mass was 9.119 kg.

Scientific focus

This reporting period falls under the ninety-eighth medium term planning period (MTP 98), which covered the period from 13 October to 9 November 2013.

Local time at ascending node (LTAN) varied from 14:40 to 17:40, thus the season was 'hot', in other words: nadir pointings would illuminate the -Z face of the spacecraft. Since quadrature continued throughout this entire reporting period, and Earth pointing for communication passes lit up a thermally sensitive face of the spacecraft, hot science observations could only be carried out if a communications pass was skipped. That allowed the spacecraft to cool down before the hot science observation was carried out.

Science priority was given to the SO2 campaign that had started in the earlier planning period MTP 95 (see status report no. 263).  To facilitate this, 2 to 3 communication passes per week were skipped. This allowed the continuation of the hot nadir pointings to make the SO2 measurements. The measurements were taken on a 2 to 4 day interval until the beginning of the atmosphere drag experiment season on 29 October. These nadir measurements of SO2 obtained with the ultraviolet channel of the SPICAV spectrometer were important for co-ordination with ground-based observers - the afternoon sector of Venus was sounded by the nadir observations, which was also the side of the planet visible from Earth.

The sixteenth occultation season started on 25 October. These Earth occultations allowed the VeRa radio science instrument to pass a downlink carrier to a ground station as the signal penetrated the Venus atmosphere. This technique sounds the atmosphere and allows temperature profiles of the atmosphere to be developed.

Change in Venus Express pericentre altitude with time from 21 July to 9 November 2013. Credit: ESA Venus Express local time at ascending node (LTAN) from 21 July to 9 November 2013. Credit: ESA
Venus Express coverage of Venus from 13 October to 9 November 2013. Credit: ESA

Payload activities
  

ASPERA The instrument was regularly operated nominally as part of the routine plan.
MAG The instrument was regularly operated nominally as part of the routine plan. MAG data that were lost due to a problem with the primary telemetry chain at Cebreros on 24 October were later retrieved, on 28 October.
PFS This instrument was not operated.
SPICAV The instrument was regularly operated nominally as part of the routine plan.
VMC The instrument was regularly operated nominally as part of the routine plan. Five single image commands did not get execution verification on 17 October. This anomaly has been observed before and is under investigation.
VeRa The instrument was regularly operated nominally as part of the routine plan.
VIRTIS The instrument was regularly operated nominally as part of the routine plan. VIRTIS data that were lost due to a problem with the primary telemetry chain at Cebreros on 24 October were later retrieved, on 28 October.

 
Future milestones

  • Aerobraking High Altitude Flight test

 
Legal disclaimer

This report is based on four ESOC mission operations reports, MOR 411 through MOR 414, as well as the MTP98 Master Science Report. Please see the copyright section of the legal disclaimer (bottom of this page) for terms of use.


Last Update: 08 April 2014

For further information please contact: SciTech.editorial@esa.int

Status Reports Archive

Images And Videos

Related Links

See Also